OpenStudy (anonymous):

TWO ANGLES ARE SUPPLEMENTARY. ONE IS 60^DEGREES MORE THAN THREE TIMES THE OTHER. FIND THE MEASURES........

OpenStudy (anonymous):

do you what the supplementary angles mean?

OpenStudy (anonymous):

do you know?

OpenStudy (anonymous):

YES THE SUM OF TWO ANGLES =180^

OpenStudy (anonymous):

RIGHT.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

....SOMEONE

OpenStudy (anonymous):

the angles are 30 & 150

OpenStudy (anonymous):

THANK YOU? CAN YOU SHOW ME HOW YOU GOT THAT OR HOW U SET THE EQUATION UP SHRIYA?

OpenStudy (anonymous):

x=3y+60 (1) x+y=180 (2) putting the value of x from (1) in (2), 3y+60+y=180 4y=120 y=30 putting in (1), x=90+60=150

OpenStudy (anonymous):

THANK YOU:)

OpenStudy (anonymous):

wc :)